# Prioritized Information Bottleneck Theoretic Framework with Distributed Online Learning for Edge Video Analytics

### v1

@@ IV-B Variational Approximation Method
The objective function of information bottleneck in Eq. (7) can be divided into two parts. The first part is $-\sum_{k=1}^{K}w_{k}\cdot I\left(Z^{(k)};Y^{(k)}\right)$, which denotes the quality of video reconstruction by decoding at an edge server. The second part is $\lambda\sum_{k=1}^{K}e^{w^{0}-w_{k}}\cdot I\left(X^{(k)};Z^{(k)}\right)$, which denotes the compression efficiency for feature extraction. As it has been shown in the way a decoder works, $p\left(Y^{\left(k\right)}|Z^{\left(k\right)}\right)$ can be any valid type of conditional distributions, but most often it is not feasible enough for straightforward calculation. Because of this complexity, it is highly challenging to directly compute the two mutual information components in Eq. (7).
As for the first part, we adopt the variational approach[50]. This approach suggests that the decoder is part of a simpler group of distributions called $Q$. We then search for a distribution $q\left(Y^{\left(k\right)}|Z^{\left(k\right)};\Theta_{d}^{(k)}\right)$ within this group that is most similar to the best possible decoder distribution, using the KL-divergence to measure the closeness. $\Theta_{d}^{(k)}$ is a learnable parameter. Because computing the high-dimensional integrals in the posterior is infeasible, we substitute the optimal inference model with a variational approximation. Thus, we obtain the lower bound of $I_{w}\left(Z^{\left(k\right)};Y^{\left(k\right)}\right)=w_{k}\cdot I\left(Z^{\left(k\right)};Y^{\left(k\right)}\right)\geq w_{k}\left\{\mathbb{E}_{p\left(Y,Z\right)}\left[\log q\left(Y^{\left(k\right)}|Z^{\left(k\right)};\Theta_{d}^{\left(k\right)}\right)\right]+H\left(Y^{\left(k\right)}\right)\right\}$, as established in Proposition 1.
@@ IV-D Network Loss Functions Derivation
$\mathcal{L}_{2}=$ $\sum_{k=1}^{K}{\underset{\mathrm{The}\ \mathrm{upper}\ \mathrm{bound}\ \mathrm{of}\ -I_{w}\left(Z^{(k)};Y^{(k)}\right)}{\underbrace{\mathbb{E}[-w_{k}\log q(Y^{(k)}|Z^{(k)};\Theta_{d}^{\left(k\right)})]}}}+\lambda\cdot\min\bigg{\{}R_{max},$ (24)
@@
The first term of $\mathcal{L}_{2}$ ignores $H(Y)$ in Ineq. (12) because it is a constant. $R_{max}$ represents the penalty for the excessive communication cost of the variation approximation $q(Z^{(k)}|V^{(k)};\Theta_{con}^{(k)})\cdot q(V^{(k)};\Theta_{l}^{(k)})$, which captures the degradation of training decoder $p(Y^{(k)}|Z^{(k)})$. In Sec. IV-C, the Multi-Frame Correlation Model leverages temporal dynamics, which is critical for sequential data processing in video analytics. The third loss function, $\mathcal{L}^{(k)}_{2}$, is needed to minimize the KL divergence between the true distribution of frame sequences and the modeled variational distribution:
@@
where $Z_{<t}^{(k)}=(Z_{t-1}^{(k)},...,Z_{t-\tau}^{(k)})$. These loss functions collectively aim to optimize the trade-off between data transmission costs and perceptual accuracy, crucial for enhancing the performance of edge analytics in multi-camera systems. Alg. 1 introduces the detailed procedure of feature extraction and variational approximation.
@@ IV-E1 Distributed Online Learning for CMAB Problem
The intuitive idea behind using distributed UCB is to manage dynamic CSI and ROI efficiently. Each camera agent independently explores and exploits available edge servers based on local observations, making the system robust to changing network conditions. The algorithm has two phases: exploration and exploitation. In the exploration phase, each agent gathers information on potential rewards. In the exploitation phase, the agent selects the best action based on the UCB value, balancing exploration and exploitation under uncertainty. The UCB value for edge server $s$ at time $t$ is $\text{UCB}_{k,{s}}(t)=\hat{\mu}_{k,{s}}(t)+\alpha\sqrt{\frac{2\ln t}{N_{k,{s}}(t)}}$, where $\hat{\mu}_{k,{s}}(t)$ is the estimated reward, driving exploitation. The second term ($\alpha\sqrt{\frac{2\ln t}{N_{k,{s}}(t)}}$) accounts for uncertainty, encouraging exploration[51]. The key intuition is that actions that have been selected fewer times carry more uncertainty, so they are given a higher bonus to encourage exploration. Conversely, as an action is selected more often and its reward estimate becomes more reliable, the bonus decreases, leading to more exploitation of that action. The parameter $\alpha$ balances how aggressively the algorithm explores uncertain actions versus exploiting known rewards. The square root component diminishes as more information is gathered, while the logarithmic term ensures the exploration bonus decreases slowly, encouraging exploration of less frequently chosen actions. Alg. 2 provides the pseudocode for the proposed gate mechanism with distributed online learning.
@@
As for distributed execution of Alg. 2, each camera agent independently updates its channel state and edge server load based on local observations (Line 7). This allows each agent to learn the optimal transmission strategy without centralized control. Moreover, each agent computes the UCB value for each edge server and selects the action that maximizes the UCB value (Lines 9-11). This decentralized decision-making process ensures scalability and adaptability in dynamic environments. To ensure consistency across agents, periodic communication rounds are initiated (Line 19). During these rounds, each agent shares its local reward estimates and action counts with a designated central edge server. This central server aggregates the information and updates the global estimates, which are then shared with all agents (Lines 19-24). This step maintains synchronization across the distributed network without requiring continuous centralized control.
@@ IV-E2 Regret Analysis
The regret analysis reflects the efficiency and adaptability of the algorithm in optimizing network performance. A lower regret bound indicates that the algorithm performs close to the optimal strategy, ensuring high inference accuracy and efficient resource utilization despite the dynamic environment and varying network conditions. In the context of a multi-camera sensing network, the choices between different cameras and edge servers are interdependent. Furthermore, the connections between different cameras and edge servers exhibit heterogeneity, with each super arm having different distribution parameters. Therefore, we consider a CMAB problem with a non-i.i.d. assumption and derive its regret upper bound in the following part.
@@
In a dynamic environment where the reward variance is denoted as $\sigma_{r}^{2}$, the cumulative regret $R(T)$ over $T$ of the distributed UCB algorithm with non-i.i.d rewards is bounded by:
$\begin{split}R(T)\leq O\Bigg{(}&\left(\sqrt{2\sigma_{r}^{2}}+2\mathcal{K}^{\text{arm}}\sqrt{\frac{2}{a_{N}}}\right)\mathcal{K}^{\text{arm}}\sqrt{T\ln T}\\ &+\frac{2\mathcal{K}^{\text{arm}}}{3}\ln T-2\mathcal{K}^{\text{arm}}\sqrt{\frac{2\ln T}{a_{N}}}\Bigg{)},\end{split}$ (28)
where $\mathcal{K}^{\text{arm}}=\sum_{k=K_{\min}}^{K_{\max}}{C_{K}^{k}S^{k+1}}$ represents the maximum number of arms in the optimal super arm, $\sigma_{r}$ is the standard deviation of the reward, and $a_{N}$ is an upper bound on the linear growth rate of the number of times UCB algorithm arms are selected over time [^5] It indicates that $N_{k}(t)$ follows a linear growth trend, i.e., $N_{k}(t)\approx a_{N}t$..
@@
It is assumed that there are $N$ camera agents and $T$ time steps. The computational complexity of the proposed DOL method in Alg. 2 for each camera agent at each time step is $O(\mathcal{K}^{\text{arm}}\log\mathcal{K}^{\text{arm}})$. The overall time complexity is $O(TN\mathcal{K}^{\text{arm}}\log\mathcal{K}^{\text{arm}})$.
@@
Assuming there are $\mathcal{X}$ communication rounds over $T$ time steps, the total communication cost of Alg. 2 is $O\left(\mathcal{X}N\mathcal{K}^{\text{arm}}\right)$.
@@ V-A Simulation Setup
Our simulations use a 2.4 GHz operating frequency, a path loss exponent of 3.5, and a shadowing deviation of 8 dB. Devices emit an interference power of 0.1 Watts, with densities ranging from 10 to 100 devices per 100 square meters, allowing us to test different levels of congestion. The bandwidth is set to 2 MHz, with cameras located at about 200 meters from the edge server. We employ the Wildtrack dataset from EPFL, which features high-resolution images from seven cameras located in a public area, capturing unscripted pedestrian movements[53]. This dataset provides 400 frames per camera at 2 frames per second, documenting over 40,000 bounding boxes that highlight individual movements across more than 300 pedestrians. Our code will be available at github.com/fangzr/PIB-Prioritized-Information-Bottleneck-Framework.
@@
Figure 7: Communication bottleneck vs MODA.
Figure 8: Communication bottleneck vs MODP.
@@
Nevertheless, the benefit of collaborative perception is accompanied by excessive communication overhead. The communication bottleneck refers to network capacity constraints that prevent real-time data transmission, causing frame latency. This issue is prevalent in UDP-based wireless streaming systems, where high throughput often results in out-of-order or delayed frames due to varying channel quality and jitter. Moreover, different coding schemes cause varying delays in dynamic channel conditions, misaligning data fusion due to channel quality and jitter. Therefore, in order to evaluate how latency differences affect perception accuracy, we set communication bottleneck constraints. In our experiments, we use MODA (Multiple Object Detection Accuracy) and MODP (Multiple Object Detection Precision) to assess coding efficiency and robustness.
In Fig. 7, PIB exhibits higher MODA across different communication bottlenecks compared to five baselines by more than 17.8%. This is due to PIB’s strategic multi-view feature fusion, informed by channel quality and priority-based ROI selection. PIB prioritizes the shared information to mitigate delays that could degrade multi-camera perception accuracy. Interestingly, JPEG outperforms video coding schemes like H.265 and AV1 in our experiments, due to the low FPS of 2 used for video transmission, which does not leverage motion prediction advantages. AV1 performs well due to its high compression efficiency compared to H.264 and H.265. Fig. 8 shows that PIB achieves higher MODP performance compared to three other baselines. The results indicate that MODP is less affected by latency because it measures the precision of detection without considering missed detections, whereas MODA is more impacted as it accounts for both missed and false detections.
Figure 9: The number of delayed cameras vs MODA.
Fig. 9 depicts the performance rates of different compression techniques in a multi-view scenario in terms of the number of delayed cameras. Our proposed PIB method and TOCOM-TEM, both utilizing multi-frame correlation models, effectively reduce redundancy across multiple frames, achieving superior MODA at equivalent compression rates. PIB, in particular, employs a prioritized IB framework, enabling an adaptive balance between compression rate and collaborative sensing accuracy, optimizing MODA across various channel conditions. It is worth noting that the impact on collaborative perception MODA can be ignored in scenarios with fewer delayed cameras (<3). However, as channel conditions worsen and more cameras experience frame delays due to failing to meet communication bottleneck constraints, the performance significantly degrades.
@@
(a) Streaming packet size for PIB, TOCOM-TEM, and AV1.
(b) Streaming packet size for JPEG, H.265, and H.264.
(c) CDF for streaming packet sizes.
Figure 11: Fig. 11(a) and Fig. 11(b) show the streaming packet sizes for various compression algorithms over time slots. Fig. 11(c) illustrates the cumulative distribution functions (CDF) of the different methods.
Fig. 11(a) shows the streaming packet size for PIB, TOCOM-TEM, and AV1 over a period of 3 minutes. For a fair comparison, all coding methods are selected at their highest MODA with the minimum communication cost data. PIB has the smallest streaming packet size among the three methods. Additionally, both PIB and TOCOM-TEM exhibit less fluctuation in packet size compared to AV1, ensuring better transmission robustness under poor channel conditions. Fig. 11(b) reflects that JPEG compression results in smaller and less fluctuating streaming packet sizes compared to H.264 and H.265. This is because the raw data transmission rate is only 2 fps, limiting the efficiency of video compression codecs. Fig. 11(c) illustrates the CDF of streaming packet sizes for all algorithms. The Standard Deviation (SD) of each algorithm is also calculated by $\text{SD}=\sqrt{\frac{\sum_{i=1}^{n}(\text{Packet Size}_{i}-\text{Mean})^{2}}{n}}$. The advantage of smaller SD is to improve transmission robustness, particularly in environments with poor channel conditions, as it leads to less jitter and requires smaller buffer sizes. As shown in Fig. 11(c), PIB has the smallest SD (3.84 KB), indicating the least fluctuation in packet sizes. This is followed by TOCOM-TEM (4.24 KB) and JPEG (6.18 KB). Moreover, five baselines exhibit higher SD values. Therefore, PIB not only significantly reduces the streaming transmission requirements but also has a smaller SD, indicating more consistent packet sizes.
TABLE I: Impact of the Number of Fusion Cameras on Collaborative Perception Accuracy and Communication Cost.
@@
Our priority-based mechanism selects the camera with the most targets within the RRoI as the highest transmission priority on the channel. Consequently, the collaboration priority is determined by the number of targets in each camera’s perception area. As shown in Table I, a mere 0.82 KB of perception data can achieve a MODA accuracy of up to 55.17% and a MODP accuracy of 69.72%. This indicates a significant redundancy among the perception data acquired from multiple edge cameras. As the number of collaborative cameras increases, the communication cost sent to the edge server also increases. Initially, with the addition of more cameras, the perception performance improves significantly. However, the performance gain from further collaboration gradually diminishes, demonstrating the principle of diminishing marginal returns.
@@
In Fig. 12, the communication bottleneck versus latency for different algorithms is shown. Although PIB increases the decoding complexity compared to TOCOM-TEM and JPEG, it significantly reduces redundant data transmission by employing a priority information bottleneck framework that selects cameras with the most prediction-relevant information. Additionally, the distributed online learning UCB algorithm filters out useless data. Consequently, PIB demonstrates higher compression efficiency compared to various video coding methods. Notably, PIB’s encoder latency is only 0.27% of that of the widely adopted AV1 compression codec.

### v2

@@ IV-B Variational Approximation Method
The objective function of information bottleneck in Eq. (7) can be divided into two parts. The first part is $-\sum_{k=1}^{K}w_{k}\cdot I\left(Z^{(k)};Y\right)$, which denotes the quality of video reconstruction by decoding at an edge server. The second part is $\lambda\sum_{k=1}^{K}e^{w^{0}-w_{k}}\cdot I\left(X^{(k)};Z^{(k)}\right)$, which denotes the compression efficiency for feature extraction. As it has been shown in the way a decoder works, $p\left(Y|Z^{\left(k\right)}\right)$ can be any valid type of conditional distributions, but most often it is not feasible enough for straightforward calculation. Because of this complexity, it is highly challenging to directly compute the two mutual information components in Eq. (7).
As for the first part, we adopt the variational approach[52]. This approach suggests that the decoder is part of a simpler group of distributions called $Q$. We then search for a distribution $q\left(Y|Z^{\left(k\right)};\Theta_{d}^{(k)}\right)$ within this group that is most similar to the best possible decoder distribution, using the KL-divergence to measure the closeness. $\Theta_{d}^{(k)}$ is a learnable parameter. Because computing the high-dimensional integrals in the posterior is infeasible, we substitute the optimal inference model with a variational approximation. Thus, we obtain the lower bound of $I_{w}\left(Z^{\left(k\right)};Y^{\left(k\right)}\right)=w_{k}\cdot I\left(Z^{\left(k\right)};Y^{\left(k\right)}\right)\geq w_{k}\left\{\mathbb{E}_{p\left(Y,Z\right)}\left[\log q\left(Y^{\left(k\right)}|Z^{\left(k\right)};\Theta_{d}^{\left(k\right)}\right)\right]+H\left(Y^{\left(k\right)}\right)\right\}$, as established in Proposition 1.
@@ IV-D Network Loss Functions Derivation
$\mathcal{L}_{2}=$ $\sum_{k=1}^{K}{\underset{\mathrm{The}\ \mathrm{upper}\ \mathrm{bound}\ \mathrm{of}\ -I_{w}\left(Z^{(k)};Y\right)}{\underbrace{\mathbb{E}[-w_{k}\log q(Y|Z^{(k)};\Theta_{d}^{\left(k\right)})]}}}+\lambda\cdot\min\bigg{\{}R_{max},$ (24)
@@
The first term of $\mathcal{L}_{2}$ ignores $H(Y)$ in Ineq. (12) because it is a constant. $R_{max}$ represents the penalty for the excessive communication cost of the variation approximation $q(Z^{(k)}|V^{(k)};\Theta_{con}^{(k)})\cdot q(V^{(k)};\Theta_{l}^{(k)})$, which captures the degradation of training decoder $p(Y|Z^{(k)})$. In Sec. IV-C, the Multi-Frame Correlation Model leverages temporal dynamics, which is critical for sequential data processing in video analytics. The third loss function, $\mathcal{L}^{(k)}_{2}$, is needed to minimize the KL divergence between the true distribution of frame sequences and the modeled variational distribution:
@@
where $Z_{<t}^{(k)}=(Z_{t-1}^{(k)},...,Z_{t-\tau}^{(k)})$. These loss functions collectively aim to optimize the trade-off between data transmission costs and perceptual accuracy, crucial for enhancing the performance of edge analytics in multi-camera systems. 1 introduces the detailed procedure of feature extraction and variational approximation.
@@ IV-E1 Distributed Online Learning for CMAB Problem
The intuitive idea behind using distributed UCB is to manage dynamic CSI and ROI efficiently. Each camera agent independently explores and exploits available edge servers based on local observations, making the system robust to changing network conditions. The algorithm has two phases: exploration and exploitation. In the exploration phase, each agent gathers information on potential rewards. In the exploitation phase, the agent selects the best action based on the UCB value, balancing exploration and exploitation under uncertainty. The UCB value for edge server $s$ at time $t$ is $\text{UCB}_{k,{s}}(t)=\hat{\mu}_{k,{s}}(t)+\alpha\sqrt{\frac{2\ln t}{N_{k,{s}}(t)}}$, where $\hat{\mu}_{k,{s}}(t)$ is the estimated reward, driving exploitation. The second term ($\alpha\sqrt{\frac{2\ln t}{N_{k,{s}}(t)}}$) accounts for uncertainty, encouraging exploration[53]. The key intuition is that actions that have been selected fewer times carry more uncertainty, so they are given a higher bonus to encourage exploration. Conversely, as an action is selected more often and its reward estimate becomes more reliable, the bonus decreases, leading to more exploitation of that action. The parameter $\alpha$ balances how aggressively the algorithm explores uncertain actions versus exploiting known rewards. The square root component diminishes as more information is gathered, while the logarithmic term ensures the exploration bonus decreases slowly, encouraging exploration of less frequently chosen actions. Algorithm 2 provides the pseudocode for the proposed gate mechanism with distributed online learning.
@@
In Algorithm 2, each camera agent independently updates its observed channel state and edge server load (Line 7). This enables decentralized learning of the optimal transmission strategy without centralized control. Specifically, in Lines 9-11, each agent computes the UCB value $\text{UCB}_{k,{s}}(t)=\hat{\mu}_{k,{s}}(t)+\alpha\sqrt{\frac{2\ln t}{N_{k,{s}}(t)}}$ for each edge server $s$, where $\hat{\mu}_{k,{s}}(t)$ is the estimated reward and $\alpha$ is used to adjust the balance between exploration and exploitation. By selecting the action with the highest UCB value, the agent optimally balances exploring the underutilized connections and exploiting the high-reward connections based on its local observations, guiding the system towards an optimal transmission strategy. Lines 15-17 handle updates to action counts, cumulative rewards, and reward estimates, refining the UCB values. Periodic communication rounds (Line 19) ensure consistency across agents by allowing each to share local reward estimates with a central server, which aggregates these data and updates global estimates for synchronization (Lines 19-24). This process maintains synchronization across the distributed network while supporting scalability and adaptability.
@@ IV-E2 Regret Analysis
The regret analysis reflects the efficiency and adaptability of the algorithm in optimizing network performance. A lower regret bound indicates that the algorithm performs close to the optimal strategy, ensuring high inference accuracy and efficient resource utilization despite the dynamic environment and varying network conditions. In the context of a multi-camera sensing network, the choices between different cameras and edge servers are interdependent. Furthermore, the connections between different cameras and edge servers exhibit heterogeneity, with each super arm having different distribution parameters. Therefore, we consider a CMAB problem with a non-identically distributed (non-i.i.d) assumption and derive its regret upper bound in the following part.
@@
In a dynamic environment with non-identically distributed (non-i.i.d) rewards due to network heterogeneity, where the variance of the reward for arm $k$ is denoted as $\sigma_{r_{k}}^{2}$, the cumulative regret $R(T)$ over $T$ rounds of the distributed UCB algorithm is bounded by:
$\begin{split}R(T)\leq O\Bigg{(}&\left(\sqrt{2\sum_{k=1}^{\mathcal{K}^{\text{arm}}}\sigma_{r_{k}}^{2}}+2\mathcal{K}^{\text{arm}}\sqrt{\frac{2}{a_{N}}}\right)\mathcal{K}^{\text{arm}}\sqrt{T\ln T}\\ &+\frac{2\mathcal{K}^{\text{arm}}}{3}\ln T-2\mathcal{K}^{\text{arm}}\sqrt{\frac{2\ln T}{a_{N}}}\Bigg{)},\end{split}$ (28)
where $\mathcal{K}^{\text{arm}}=\sum_{k=K_{\min}}^{K_{\max}}{C_{K}^{k}S^{k+1}}$ represents the maximum number of arms in the optimal super arm, $\sigma_{r_{k}}$ is the standard deviation of the reward for arm $k$, and $a_{N}$ is an upper bound on the linear growth rate of the number of times UCB algorithm arms are selected over time [^6] It indicates that $N_{k}(t)$ follows a linear growth trend, i.e., $N_{k}(t)\approx a_{N}t$.. The non-i.i.d nature of the rewards represents the heterogeneity in the network where different arms can have different reward distributions due to varying network conditions, processing capabilities, and data qualities.
@@
It is assumed that there are $N$ camera agents and $T$ time steps. The computational complexity of the proposed DOL method in Algorithm 2 for each camera agent at each time step is $O(\mathcal{K}^{\text{arm}}\log\mathcal{K}^{\text{arm}})$. The overall time complexity is $O(TN\mathcal{K}^{\text{arm}}\log\mathcal{K}^{\text{arm}})$.
@@
Assuming there are $\mathcal{X}$ communication rounds over $T$ time steps, the total communication cost of Algorithm 2 is $O\left(\mathcal{X}N\mathcal{K}^{\text{arm}}\right)$.
@@ V-A Simulation Setup
Our simulations use a 2.4 GHz operating frequency, a path loss exponent of 3.5, and a shadowing deviation of 8 dB. Devices emit an interference power of 0.1 Watts, with densities ranging from 10 to 100 devices per 100 square meters, allowing us to test different levels of congestion. The bandwidth is set to 2 MHz, with cameras located at about 200 meters from the edge server. We employ the Wildtrack dataset from EPFL, which features high-resolution images from seven cameras located in a public area, capturing unscripted pedestrian movements[55]. This dataset provides 400 frames per camera at 2 frames per second, documenting over 40,000 bounding boxes that highlight individual movements across more than 300 pedestrians. As shown in Fig. 7, our experimental setup features a practical hardware testbed that includes three distinct edge devices: NVIDIA Jetson™ Orin Nano™ 4GB, NVIDIA Jetson™ Orin NX™ 16GB, and ThinkStation™ P360. The edge devices collaboratively interact with edge servers equipped with RTX 5000 Ada GPUs for efficient video decoding. Our code will be made available at github.com/fangzr/PIB-Prioritized-Information-Bottleneck-Framework.
@@
(a) Communication bottleneck vs MODA.
(b) Communication bottleneck vs MODP.
(c) Number of delayed cameras vs MODA.
Figure 8: Impact of communication bottlenecks and delayed cameras on perception accuracy.
@@
To evaluate the impact of communication bottlenecks and delayed cameras on perception accuracy, we present in Fig. 8(a)8(c) the relationships between communication constraints and the perception accuracy. Nevertheless, the benefit of collaborative perception is accompanied by excessive communication overhead. The communication bottleneck refers to network capacity constraints that prevent real-time data transmission, causing frame latency. This issue is prevalent in UDP-based wireless streaming systems, where high throughput often results in out-of-order or delayed frames due to varying channel quality and jitter. Moreover, different coding schemes cause varying delays in dynamic channel conditions, misaligning data fusion due to channel quality and jitter. Therefore, in order to evaluate how latency differences affect perception accuracy, we set communication bottleneck constraints. In our experiments, we use MODA (Multiple Object Detection Accuracy) and MODP (Multiple Object Detection Precision) to assess coding efficiency and robustness.
In Fig. 8(a), PIB exhibits higher MODA across different communication bottlenecks compared to five baselines by more than 17.8%. This is due to PIB’s strategic multi-view feature fusion, informed by channel quality and priority-based ROI selection. PIB prioritizes the shared information to mitigate delays that could degrade multi-camera perception accuracy. Interestingly, JPEG outperforms video coding schemes like H.265 and AV1 in our experiments, due to the low FPS of 2 used for video transmission, which does not leverage motion prediction advantages. AV1 performs well due to its high compression efficiency compared to H.264 and H.265. Fig. 8(b) shows that PIB achieves higher MODP performance compared to three other baselines. The results indicate that MODP is less affected by latency because it measures the precision of detection without considering missed detections, whereas MODA is more impacted as it accounts for both missed and false detections.
Fig. 8(c) depicts the performance rates of different compression techniques in a multi-view scenario in terms of the number of delayed cameras. Our proposed PIB method and TOCOM-TEM, both utilizing multi-frame correlation models, effectively reduce redundancy across multiple frames, achieving superior MODA at equivalent compression rates. PIB, in particular, employs a prioritized IB framework, enabling an adaptive balance between compression rate and collaborative sensing accuracy, optimizing MODA across various channel conditions. It is worth noting that the impact on collaborative perception MODA can be ignored in scenarios with fewer delayed cameras (<3). However, as channel conditions worsen and more cameras experience frame delays due to failing to meet communication bottleneck constraints, the performance significantly degrades.
@@
(a) Streaming packet size for PIB and five baselines over time.
(b) CDF for streaming packet sizes.
Figure 10: (a) Streaming packet sizes for various compression algorithms over time slots. (b) Cumulative Distribution Functions (CDF) of the streaming packet sizes for different methods.
Fig. 10 presents the streaming packet sizes and their cumulative distribution functions (CDF) for various compression algorithms. Fig. 10(a) illustrates the streaming packet sizes for PIB, TOCOM-TEM, AV1, JPEG, H.265, and H.264 over a duration of three minutes. All encoding methods were evaluated at their highest MODA with minimal communication costs. PIB consistently exhibits the smallest packet sizes, followed by TOCOM-TEM, indicating superior transmission efficiency. Additionally, PIB and TOCOM-TEM demonstrate less variability in packet sizes compared to AV1, enhancing transmission robustness under adverse channel conditions. JPEG compression yields smaller and more stable packet sizes than H.264 and H.265, likely due to the limited transmission rate of 2 fps restricting the efficiency of video codecs. Fig. 10(b) shows the CDF of streaming packet sizes for all algorithms. The standard deviation (SD) for each method is calculated as $\text{SD}=\sqrt{\frac{\sum_{i=1}^{n}(\text{Packet Size}_{i}-\text{Mean})^{2}}{n}}$. A lower SD indicates improved transmission robustness by reducing jitter and minimizing buffer requirements. PIB has the lowest SD (3.84 KB), followed by TOCOM-TEM (4.24 KB) and JPEG (6.18 KB). The other baseline methods exhibit higher SD values, underscoring PIB’s advantage in minimizing both transmission requirements and packet size variability.
TABLE II: Impact of the Number of Fusion Cameras on Collaborative Perception Accuracy and Communication Cost.
@@
TABLE III: Encoder Latency Across Different Platforms.
Phase Platform Nano (ms) Orin NX (ms) P360 (ms)
Feature map generation 755.32±69.32 227.54±2.65 37.49±0.90
Entropy coding 10.83±3.51 1.79±0.75 0.40±0.11
Total encoder latency 766.15±70.55 229.34±2.67 37.80±0.94
Our priority-based mechanism selects the camera with the most targets within RoI for the highest transmission priority. Collaboration priority is thus determined by the target count in each camera’s perception area. As shown in Table II, just 0.82 KB of perception data achieves a MODA accuracy of 55.17% and MODP of 69.72%, highlighting significant redundancy among edge cameras. Adding more cameras initially improves perception significantly but offers diminishing returns as communication costs to the edge server increase. In Fig. 11, we show the relationship between communication bottleneck and total latency for different algorithms. By leveraging a priority information bottleneck framework and the UCB algorithm to reduce redundancy, our PIB, despite slightly higher encoding latency than the traditional video codecs, can achieve much lower transmission latency due to its efficient compression. Under a 25 KB/s bottleneck, our PIB reduces latency by 25.1% over TOCOM-TEM and 74.8% over JPEG. At 50 KB/s, our PIB outperforms TOCOM-TEM by 18.3% and JPEG by 65.1%, respectively. At 100 KB/s, PIB achieves 10.8% and 46.1% lower latency than TOCOM-TEM and JPEG, respectively. The encoding latency results of our PIB in different edge devices are presented in Table III. It can be observed that the feature map generation phase dominates the overall encoding latency, while the entropy coding phase contributes a negligible amount of time. Furthermore, edge devices with higher computating capacity exhibit significantly lower encoding latency.